OpenStudy (djcall11):

I don't know how to begin A boy is riding a bicycle at a velocity of 5.0 m/s. He applies the brakes and uniformly negatively accelerates to a stop at a rate of 2.5 m/s^2. How long does it take for the bicycle to stop?

OpenStudy (djcall11):

I this it's just a simple answer 2.0 s

OpenStudy (danjs):

If you are starting at an initial velocity, and have a uniform acceleration, you can use Change in velocity = acceleration * time accelerating

OpenStudy (danjs):

\[\large V _{2}-V _{1}=a*t\] Your final velocity V2, starting at Velocity V1, and constantly accelerating at a, for time t. \[\large V _{2}=V _{1}+a*t\]

OpenStudy (djcall11):

2.5=5*0.5 2.5=2.5

OpenStudy (danjs):

V1 = 5.0 m/s V2 = 0.0 m/s a = -2.5 m/s^2 0 = 5.0 + (-2.5)*t

OpenStudy (danjs):

2.5*t = 5.0 t = 2.0 s

OpenStudy (djcall11):

thank you